OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x^8 3x^6) cubic feet. What polynomial represents the height of the box in feet? A. 4x 3 B. 4x^6 3x^4 C. 4x^8 3x^6 D. 4x^7 3x^5

OpenStudy (raden):

v = lwh

OpenStudy (anonymous):

#11. x = length of box, x= width of box. height of box = ? Volume of box = length * width * height = 4x^8 + 3x^6. Volume of box = length * width * height = x * x * height x * x * height = 4x^8 + 3x^6 x^2 * height = 4x^8 + 3x^6 x^2 * height / x^2 = (4x^8 + 3x^6) / x^2 height = x^2 * (4x^6 + 3x^4) / x^2 height = 4x^6 + 3x^4

OpenStudy (anonymous):

i stilll dont understand

OpenStudy (anonymous):

i have another question charlene made the sketch below in order to find the height x of a pole. She positioned a mirror on the ground so that she could see the reflection of the top of the pole.Her height,her distance from the mirror, and her line of sight to the mirror determine the smallest triangle.The poles height, its distance from the mirror,and the distance from the top of the pole to the mirror form a larger similar triangle. Find the height of the pole to the nearest tenth. a.16.5ft b.22.ft c.20.5ft d.17.5ft